# Scalable Hybrid quantum Monte Carlo simulation of U(1) gauge field coupled to fermions on GPU

### v1

@@ II.7 Complexity analysis
As shown in Fig. 7, both the sample generation and stochastic estimation display nearly linear scaling behavior with respect to $N_{\tau}V_{s}$ as expected, so does the overall time complexity. Moreover, the stochastic estimation has slightly larger scaling power $0.905$ than sample generation $1.313$, and also larger latency at large system sizes. This is because stochastic estimation needs to simultaneously process $f(N_{\textmd{rv}})$ input vectors, while sample generation only needs to process one. Thus, the stochastic estimation has much stronger arithmetic intensity, and the latency scales faster with respect to system size. Nevertheless, Fig. 7 also shows a green dashed guideline which is the computational complexity of DQMC. At system size of $L^{3}\sim 10^{5}$, the two algorithms already have two orders of magnitude latency difference, which shows the advantage of the HQMC.
III Results
Here we show the results from our GPU-accelerated HQMC simulations, we focuse on the non-compact gauge field case, and will present results on the compact case in SM [64]. For small system sizes, we benchmark the HQMC with the DQMC and find consistency at the quantitative level, up to sizes of $160\times 16\times 16$ ($N_{\tau}\times L\times L$). We then use HQMC to push for the thermodynamic limit, up to $660\times 66\times 66$, which is more than $50$ times larger in the space-time volume limit of DQMC, and the data clearly reveal the power-law correlations in different fermion bilinear operators and the flux operator inside the $U(1)$ DSL phase.
III.1 Physical observables
We first introduce the fermionic observables used to characterize and detect the $U(1)$ DSL phase. In our problem of QED3 $U(1)$ gauge-field coupled fermions, one caution we need to take is the gauge invariance of an observable. For example, the single-particle Green’s function is not gauge-invariant, and to make it gauge-invariant, one will require an extra integral of the $U(1)$ phase along the path connecting the two fermions. On the other hand, the fermion bilinear operators are gauge-invariant, and we measure the following two correlation functions of fermion bilinears – spin-spin and bond-bond correlation functions – as
$C_{\textmd{B}}(i,j)$ $=\langle B_{i}B_{j}\rangle-\langle B_{i}\rangle\langle B_{j}\rangle,$ (30)
where the spin operator is defined as $S^{\alpha\beta}(i)=\psi_{i\alpha}^{\dagger}\psi_{i\beta}-\frac{1}{N_{f}}\delta_{\alpha\beta}\sum_{\gamma}\psi_{i\gamma}^{\dagger}\psi_{i\gamma}$ , bond operator along the nearest-neighbor bond in $\hat{x}$ direction is defined as $B_{i}=\sum_{\alpha}(\psi^{\dagger}_{i,\alpha}e^{i\phi_{i,i+\hat{x}}}\psi_{i+\hat{x},\alpha}+\text{h.c.})$, and $\alpha,\beta\in\left\{\uparrow,\downarrow\right\}$ are spin flavor indices. Here, the spin operators are written as antisymmetric self-conjugate representation [1], and are related to the normal spin operators definition as $S^{\uparrow\downarrow}=S^{+},S^{\downarrow\uparrow}=S^{-},S^{\uparrow\uparrow}=S^{z},S^{\downarrow\downarrow}=-S^{z}$. The corresponding structure factors are defined as the Fourier transformations of the correlation functions $\chi_{\textmd{S, B}}(\mathbf{k})=\frac{1}{V_{s}^{2}}\sum_{ij}C_{\textmd{S, B}}(i,j)e^{i\mathbf{k}\cdot(\mathbf{r}_{i}-\mathbf{r}_{j})}$. The spin form factor $\chi_{S}(\mathbf{k})$ evaluated at $X=(\pi,\pi)$ point will be used to characterize the antiferromagnet (AFM) order, while the bond structure factor $\chi_{B}(\mathbf{k})$ evaluated at $M=(\pi,0)$ point, will be used to characterize the columnar VBS order.
Other than the gauge invariant fermionic observables, we also measure the correlation of the magnetic flux operator $\theta_{\square}(\tau)$, with a focus on its imaginary time dynamics. Since, in the static limit ($J=0$) of the QED3, the ground energy magnetic configuration is known to be $\pi$-flux on a square lattice [61]. When $J$ is turned on, the fluctuation of magnetic flux is centered also around $\pi$-flux. So, to study the correlation of the magnetic fluxes, we measure the correlation of its sine value instead, in order to have a clean background, i.e. $\langle\sin\theta_{\square}\rangle=0$. We define
III.2 The operator scaling dimensions in non-compact QED3
As mentioned in the Introduction (Sec. I.1), the DSL with non-compact QED3 has been analyzed extensively and is believed to describe a scale-invariant conformal field theory (CFT). Consequently, power law behavior is expected for all gauge-invariant observables. At our setting of $N_{f}=2$, the CFT is expected to have emergent $SU(4)$ symmetry (with two cones in the BZ), one expects the scaling dimension of the fermion bilinears or mass terms: a singlet $\Delta_{s}\approx 2.3$ and a set of adjoints $\Delta_{\text{adj}}\in(1.4,1.7)$, where there are 15 of them ranging over the $SU(4)$ generators. These scaling dimensions are obtained from the $\epsilon$-expansion from $d=4-2\epsilon$ to 3 [18]. We note that, the spin operator and the bond operator mentioned above in Eqs. (29) and (30) are among the $15$ adjoint fermion bilinear terms and consequently they are expected to share the same power-law ($\sim 1/r^{2\Delta_{\text{adj}}}$) when we compute their correlation functions [3, 6] and it is the indeed the case as we show in the next section.
Moreover, in the gauge sector, one expects that the $U(1)$ DSL should exhibit flux-flux correlations dictated by $\Delta_{J}=2$, which implies the long-time correlations of the local flux should behave as $\langle\sin\theta_{\Box}(\tau)\sin\theta_{\Box}(0)\rangle\sim\tau^{-4}$. In contrast, the flux correlation in a non-compact $U(1)$ gauge field theory has a power-law dynamic correlation of $\langle\sin\theta_{\Box}(\tau)\sin\theta_{\Box}(0)\rangle\sim\tau^{-3}$, which is non-conformal but instead from a Goldstone mode of spontaneously broken magnetic $U(1)$ symmetry, related to flux conservation [82]. From our HQMC simulation, the $\tau^{-4}$ behavior of flux correlation is seen at large space-time volume.
Figure 8: The spin correlation function for non-compact QED3. The log-log plot of the spin-spin correlation $C_{S}^{\uparrow\downarrow}(r,0)$ as a function of distance $r$ for non-compact QED3in Eqs. (1) and (3) with $K=0$ and $J=1.25$ , for various cubic lattice sizes shown as $N_{\tau}\times L^{2}$ in the legend. Notice that $C_{S}^{\uparrow\downarrow}=C_{S}^{\downarrow\uparrow}$, so we only need to measure $C_{S}^{\uparrow\downarrow}$ here. To avoid the even-odd oscillation in the finite-size data, we only plot the data of the distance $r=\textmd{odd}$ points. The gray square, triangular and diamond dots show the results computed from the DQMC to benchmark the HQMC result of the corresponding lattice sizes. At the large distance and the large system sizes, the data are manually fitted by the black solid line, which displays a power law decay as $1/r^{3.0}$.
Figure 9: The bond correlation function for non-compact QED3. The log-log plot of the bond-bond correlation (along $\hat{x}$ direction) as a function of distance $r$ for non-compact QED3 in Eqs. (1) and (3) with $K=0$ and $J=1.25$ , for various lattice sizes shown as $N_{\tau}\times L^{2}$ in the legend. At the large distance and the large system sizes, the data are manually fitted by the black line, which displays a power law decay as $1/r^{3.3}$.
Figure 10: The flux correlation function for non-compact QED3. The log-log plot of the flux-flux correlation as a function of imaginary time $\tau$ for non-compact QED3 with $K=0$ and $J=1.25$ Eqs. (1) and (3), for various lattice sizes, shown as $N_{\tau}\times L^{2}$ in the legend. At the large distance and the large system sizes, the data are manually fitted by the black line, which displays a power law decay as $\tau^{-4}$. The black dashed line shows the $\tau^{-3}$ power law behavior for comparison.
III.3 Numerical results
In this section, we analyze the numerical results of the fermionic correlation functions and the gauge flux correlation functions in non-compact QED3 in Eqs. (1) and (3) as introduced above, with a focus on their scaling dimensions and how they shed light on the existence of the $U(1)$ DSL phase.
Fig. 8 shows the log plot of the spin-spin correlation $C_{S}^{\uparrow\downarrow}(r,0)$ as a function of real space distance $r$ at $J=1.25,K=0$, for various lattice sizes. As the system size increases from $L=12$ all the way to $L=60$, the spin correlation curves monotonically unroll, and approach a straight envelope line at large distances. This indicates a power law decay behavior. By manually fitting this envelop curve shown as the dark solid line, we obtain the exponent of the power law decay to be around $3.0$. This agrees with the scaling dimension of the fermion bilinear operators $2\Delta_{\textmd{adj}}\in(2.8,3.4)$ predicted by the CFT, where the fermion bilinear operators serve as the adjoint representation of the emergent $SU(4)$ symmetry. We note that around $10^{-5}$ in Fig. 8, the data fall below the noise level. So the manual fitting is performed mainly with the few large-distance data $\gtrsim 10^{-5}$ at large system sizes. In Fig. 8, we also plot the DQMC results for several small lattice sizes to benchmark the HQMC results, and they well agree with each other. These results are complemented by the benchmark data in Sec.II of the Supplemental Materials (SM) [64].
Fig. 9 shows log plot of the bond-bond correlation $C_{B}(r,0)$ as a function of real space distance $r$ at $J=1.25,K=0$, for various lattice sizes. The data of the bond correlation generally have a larger variance than the spin correlation function shown in Fig. 8, which is due to the explicit gauge field dependence in the bond operator $B_{i}$. But such variance from gauge field fluctuation seems to be alleviated at large system sizes, and the noise level decreases to around $5\times 10^{-5}$ as the system size increases to $N_{\tau}\times L^{2}=600\times 60^{2}$. Furthermore, for system sizes $L\gtrsim 40$, the decaying behavior of the bond-bond correlation is stable around a straight line, which again indicates a power law decay. The manual linear fitting shown as the dark solid line yields a power of around $3.3$, which again agrees with the scaling dimension of the fermion bilinear operators $2\Delta_{\textmd{adj}}\in(2.8,3.4)$.
As analyzed in the above section and shown in Fig. 8 and Fig. 9, the fact that both spin and bond correlation functions display power law decay with powers consistent with the scaling dimension of the CFT fermion bilinear operators supports the existence of the $U(1)$ DSL phase in non-compact QED3 at $J=1.25$. This also agrees with a recent DQMC study on the same model [82].
Fig. 10 shows the log plot of the magnetic flux correlation function $C_{\textmd{flux}}(\tau)$ as a function of imaginary time $\tau$ for various lattice sizes, at $J=1.25,K=0$. It can be seen that, as system size increases, the curves become increasingly linear before entering the noise level at around $2\times 10^{-3}$. Such linear tendency is manually fitted with the dark solid line, which yields a power of around $4.0$. In contrast, the black dashed line shows the $\tau^{-3}$ behavior, which originates from the non-conformal flux-flux correlation in a $U(1)$ gauge field theory [83]. This is clearly distinct from the $\tau^{-4}$ behavior originating from the conserved current’s scaling dimension $\Delta_{J}=2$, as mentioned above. Thus, the power law decay in the flux correlation function with power $4.0$ again indicates the existence of $U(1)$ DSL phase in $J=1.25$ in the non-compact QED3.
IV Discussions
In this work, we developed a scalable GPU-accelerated hybrid quantum Monte Carlo (HQMC) algorithm for simulating $U(1)$ gauge field coupled to fermions in (2+1)d. The technical advances achieved here are threefold:
  1. 1.
    the construction of a problem-specific preconditioner that enables rapid convergence of the conjugate gradient solver;
  2. 2.
    the design of customized CUDA kernels that exploit the sparsity and structure of the matrix-vector operations intrinsic to the model;
  3. 3.
    the integration of CUDA graph acceleration to minimize kernel launch overhead and maximize GPU utilization.
These optimizations collectively allow us to reach a linear scaling in computational complexity with respect to space-time volume, i.e. $O(N_{\tau}V_{s})$, and to simulate unprecedentedly large system sizes with ease.
Since the HQMC gives the full space-time correlation of the fermion bilinear operators, one immediate future direction is to investigate the spectral information inside the $U(1)$ DSL phase from the stochastic analytic continuation of the HQMC data [82]. As we have both large $N_{\tau}$ and $V_{s}$, the momentum and energy resolution of the obtained spectra (for example the magnetic spectra from spin-spin correlation function) will be much higher than those from the DQMC [82]. In this way, the weight distribution and the momentum dependence of the spinon continuum spectra of $U(1)$ DSL can be revealed, and hopefully connection with experiment results on triangular lattice antiferromagnets YbZn2GaO5 [35, 36] and the $A$-YbSe2 delafossites [37, 38] and kagomé antiferromagnet YCu3(OD)6Br2[Br1-x(OD)x[39, 40, 41, 42, 43, 44, 45] can be made.
Moreover, by adding magnetic field to the QED3 model, one can investigate how the flavor chemical potential will change the $U(1)$ DSL phase and understand the emergent gauge flux in magnetized Dirac spin liquids. For example, recent DQMC work [82] and analytic works [84, 85] have explored the various symmetry-breaking phases out of $U(1)$ DSL under magnetic field. In particular, a novel chiral flux phase, where the system spontaneously generates a net emergent orbital magnetic flux (deviated from the $\pi$-flux at zero field case) has been seen from smaller size DQMC simulation. This spontaneous orbital magnetic flux changes the behavior of the spinons, and make them form the relativistic Landau levels — energy levels normally seen in electrons moving in a magnetic field. Remarkably, the system also exhibits a gapless photon-like mode, a signature of the coherent emergent gauge field, which can be potentially detected in inelastic neutron scattering experiments [82]. Our GPU-accelerated HQMC algorithm can provide systematic results of such interesting phase and rich diagram, as a function of magnetic field, gauge field fluctuations, as well as the temperature axes. One possible direction is to investigate the effect of different filling of the spinon Landau levels in the magnetized DSL phase [86], and establish possible connections to experimentally observed unusual magnetic oscillations in the YCu3(OD)6Br2[Br1-x(OD)x] kagomé antiferromagnet [45].
Acknowledgements.
@@
We thank U.F.P. Seifert, O.A. Starykh and L. Balents for the stimulating discussion on the physics of $U(1)$ DSL and QED3, and the collaboration on a related work [82]. We thank M. Ulybyshev and F. Assaad for inspiring communications on hybrid Monte Carlo [87]. KXF thanks Natalia Perkins for the collaboration in a related project. KXF, CC and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. AoE/P-701/20, 17309822, HKU C7037-22GF, 17302223, 17301924), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A HKU703/22). We thank HPC2021 system under the Information Technology Services and the Blackbody HPC system at the Department of Physics, University of Hong Kong, as well as the Beijng PARATERA Tech CO.,Ltd [88] for providing HPC resources that have contributed to the research results reported within this paper.
@@
In this supplemental material, we provide the technical details concerning the QMC measurements of conserved currents (Sec. SII) and the finite size scaling of the order parameters at the easy-plane DQCP (Sec. SIII).
@@
Figure S5: The log plot of the spin-spin correlation as a function of distance $r$ for compact QED3with $K=1$, at $J=1.25$, for various lattice sizes shown as $N_{\tau}\times L^{2}$ in the legend. To avoid the even-odd oscillation in the finite-size data, we only plot the data of the distance $r=\textmd{odd}$ points. The gray square, triangular and diamond dots show the results computed from the DQMC to benchmark the HQMC result of the corresponding lattice sizes. At the large distance and the large system sizes, the data are manually fitted by the black line, which displays a power law decay with scaling dimension of around $3.2$.
Figure S6: The log-log plot of the bond-bond correlation (along $\hat{x}$ direction) as a function of distance $r$ for compact QED3with $K=0$, at $J=1.25$, for various lattice sizes shown as $N_{\tau}\times L^{2}$ in the legend. Notice that $C_{S}^{\uparrow\downarrow}=C_{S}^{\downarrow\uparrow}$, so we only need to measure $C_{S}^{\uparrow\downarrow}$ here. To avoid the even-odd oscillation in the finite-size data, we only plot the data of the distance $r=\textmd{odd}$ points. At the large distance and the large system sizes, the data are manually fitted by the black line, which displays a power law decay with scaling dimension of around $3.2$.
Figure S7: The log-log plot of the flux-flux correlation as a function of imaginary time $\tau$ for compact QED3  with $K=1$ and $J=1.25$, for various lattice sizes, shown as $N_{\tau}\times L^{2}$ in the legend. At the large distance and the large system sizes, the data are manually fitted by the black line, which displays a power law decay with scaling dimension of around $4.1$. As lattice sizes increases, the slope still shows finite-size oscillation behavior up to size $N_{\tau}\times L^{2}=660\times 66^{2}$.
@@
Fig. S5 shows the log plot of the spin-spin correlation as a function of lattice distance $r$ for compact QED3 at $J=1.25$, for various lattice sizes. The noise level is at around $7\times 10^{-6}$. As lattice size increases, the log plot of the spin-spin correlation function consistently displays a linear decay behavior at large distance $r$, until it reaches the noise level. This agrees with the power law decay behavior of the fermionic observable, which manifests the property of the $U(1)$ DSL critical phase. The linear decay is manually fitted by the black solid line which gives a power of around $3.2$. This agrees with the scaling dimension of fermion bilinear operators $2\Delta_{\rm adj}\in(2.8,3.4)$ predicted by the CFT, where the fermion bilinear operators serve as the adjoint representation of the emergent $SU(4)$ symmetry, as introduced in the main text and Refs. [3, 18]. In Fig. S5, we also plot the DQMC results for several small lattice sizes to benchmark the HQMC results, and they agree well with each other.
Similarly, in Fig. S6 which shows the log plot of the bond-bond correlation as a function of distance $r$, the data points for various lattice sizes also consistently converge to a straight line, until they reach a noise level of $2\times 10^{-5}$. The points are large distance $r$ are manually fitted with the black solid line, which gives the power of around $3.2$. This again agrees with the scaling dimension of fermion bilinear operators $2\Delta_{\rm adj}\in(2.8,3.4)$ predicted by the CFT.
Fig. S7 shows the log plot of the flux-flux correlation as a function of the imaginary time $\tau$ for compact QED3 model with $K=1$ and $J=1.25$. As lattice size increases, the flux-flux correlation at large $\tau$ consistently converge to a clear linear behavior before they reach a noise level of around $2\times 10^{-3}$. Such limit behavior is again manually fitted by the black solid line, which gives a power of around $4.1$, and is distinct from the dashed line with the power of $3.0$. This verifies the fact that the scaling dimension of the conserved current in $U(1)$ DSL phase is $\Delta_{J}=2$.
@@
In this problem, the Green’s function of the two fermion flavors are equal $G_{\uparrow}=G_{\downarrow}$. So we will omit the flavor index in the Green’s functions next. Similarly, the bond correlation function is

### v3

@@ III.0.3 Speedup of customized CUDA kernel
As shown in Fig. 2(c), both the sample generation and stochastic estimation display nearly linear scaling behavior with respect to $N_{\tau}V_{s}$ as expected, so does the overall time complexity. Moreover, the stochastic estimation has slightly larger scaling power $0.905$ than sample generation $1.313$, and also larger latency at large system sizes. This is because stochastic estimation needs to simultaneously process $f(N_{\textmd{rv}})$ input vectors, while sample generation only needs to process one. Thus, the stochastic estimation has much stronger arithmetic intensity, and the latency scales faster with respect to system size. Nevertheless, Fig. 2(c) also shows a green dashed guideline which is the computational complexity of DQMC. At system size of $L^{3}\sim 10^{5}$, the two algorithms already have two orders of magnitude latency difference, which shows the advantage of the HQMC.
Acknowledgments
@@
We thank U.F.P. Seifert, O.A. Starykh and L. Balents for the stimulating discussion on the physics of $U(1)$ DSL and QED3, and the collaboration on a related work [chenEmergent2025]. We thank M. Ulybyshev and F. Assaad for inspiring communications on hybrid Monte Carlo. KXF thanks Natalia Perkins for the collaboration in a related project. KXF, CC and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. AoE/P-701/20, 17309822, HKU C7037-22GF, 17302223, 17301924), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A HKU703/22). We thank HPC2021 system under the Information Technology Services and the Blackbody HPC system at the Department of Physics, University of Hong Kong, as well as the Beijing Paratera Tech Corp., Ltd for providing HPC resources that have contributed to the research results reported within this paper.
@@
In this supplemental material, we provide the technical details concerning the QMC measurements of conserved currents (Sec. SII) and the finite size scaling of the order parameters at the easy-plane DQCP (Sec. SIII), and the derivation of the fermionic correlation functions (Sec. SIV).
@@
Figure S5: The correlation functions for compact QED3 with $K=1$ and $J=1.25$. (a) shows the log plot of the spin-spin correlation $C_{S}^{\uparrow\downarrow}(r,0)$ as a function of distance $r$, for various cubic lattice sizes shown as $N_{\tau}\times L^{2}$ in the legend. To avoid the even-odd oscillation in the finite-size data, we only plot the data of the distance $r=\textmd{odd}$ points. The gray square, triangular and diamond dots show the results computed from the DQMC to benchmark the HQMC result of the corresponding lattice sizes. At the large distance and the large system sizes, the data are manually fitted by the black line, which displays a power law decay with scaling dimension of around $3.9$. (b) shows the log plot of the bond-bond correlation (along $\hat{x}$ direction) as a function of distance $r$. The legend is the same as that in (a). At the large distance and the large system sizes, the manually fit line of the bond-bond correlation also displays a power law decay as $1/r^{3.7}$. (c) shows the log plot of the flux-flux correlation as a function of imaginary time $\tau$. At the large distance and the large system sizes, the black solid line shows the manual fitting which displays a power law decay as $\tau^{-4.1}$. The black dashed line shows the $\tau^{-3.0}$ power law behavior for comparison.
@@
Fig. S5(a) shows the log plot of the spin-spin correlation as a function of lattice distance $r$ for compact QED3 at $J=1.25$, for various lattice sizes. The noise level is at around $7\times 10^{-6}$. As lattice size increases, the log plot of the spin-spin correlation function consistently displays a linear decay behavior at large distance $r$, until it reaches the noise level. This agrees with the power law decay behavior of the fermionic observable, which manifests the property of the $U(1)$ DSL critical phase. The linear decay is manually fitted by the black solid line which gives a power of around $3.9$. This agrees with the scaling dimension of fermion bilinear operators $2\Delta_{\rm adj}\in(3.5,3.9)$ predicted by the CFT, where the fermion bilinear operators serve as the adjoint representation of the emergent $SU(4)$ symmetry, as introduced in the main text and Refs. [Hermele2005, di2017scaling]. In Fig. S5(a), we also plot the DQMC results for several small lattice sizes to benchmark the HQMC results, and they agree well with each other.
Similarly, in Fig. S5(b) which shows the log plot of the bond-bond correlation as a function of distance $r$, the data points for various lattice sizes also consistently converge to a straight line, until they reach a noise level of $2\times 10^{-5}$. The points are large distance $r$ are manually fitted with the black solid line, which gives the power of around $3.7$. This again agrees with the scaling dimension of fermion bilinear operators $2\Delta_{\rm adj}\in(3.5,3.9)$ predicted by the CFT.
Fig. S5(c) shows the log plot of the flux-flux correlation as a function of the imaginary time $\tau$ for compact QED3 model with $K=1$ and $J=1.25$. As lattice size increases, the flux-flux correlation at large $\tau$ consistently converge to a clear linear behavior before they reach a noise level of around $2\times 10^{-3}$. Such limit behavior is again manually fitted by the black solid line, which gives a power of around $4.1$, and is distinct from the dashed line with the power of $3.0$. This verifies the fact that the scaling dimension of the conserved current in $U(1)$ DSL phase is $\Delta_{J}=2$.
@@
In this problem, the Green’s function of the two fermion flavors are equal $G_{\uparrow}=G_{\downarrow}$. So we will omit the flavor index in the Green’s functions below. Similarly, the bond correlation function is